OpenStudy (anonymous):

Evaluate the expression. r = <7, -3, -7>, v = <4, 6, -5>, w = <-5, -6, -3> v ⋅ w -41 <-20, -36, 15> <-28, 18, -35> -9

OpenStudy (anonymous):

v ⋅ w=v1w1+v2w2+v3w3=-20-36+15=-41